use the discriminant to determine the number of solutions of the quadratic equation, and whether the solutions are real or complex. Note: It is not necessary to find the roots; just determine the number and types of solutions.
x2 + 6x - 7 = 0
z2 + z + 1 = 0
(3)1/2y2 - 4y - 7(3)1/2 = 0
2x2 - 10x + 25 = 0
2x2 - 6x + 5 = 0

Answer by venugopalramana(3286)

use the discriminant to determine the number of solutions of the quadratic equation, and whether the solutions are real or complex. Note: It is not necessary to find the roots; just determine the number and types of solutions.
x2 + 6x - 7 = 0
DEL=6^2-4*1*(-7)=36+28=64..2 SOLUTIONS,REAL
z2 + z + 1 = 0
DEL=1-4=-3..-VE........IMAGINARY SOLUTION ..2
(3)1/2y2 - 4y - 7(3)1/2 = 0...????PLEASE TYPE PROPERLY
2x2 - 10x + 25 = 0
DEL=100-4*2*25=-VE......IMAGINARY..2
2x2 - 6x + 5 = 0
DEL=36-40=-VE...IMAGINARY..2